### v2

@@ 4.2 Dataset Preparation and Calibration
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 1: An illustration of the performance of a CIFAR-10 classifier. Each graph corresponds to a single output category. The $Y$-axes in the graphs show the average number of hits (bold line) and standard deviation (faded area), for $k=1000$ perturbations around genuine (blue) and adversarial (orange) images; whereas the $X$-axes represent different values of $\epsilon$ (in percents). The goal is to have as little overlap as possible between the two shaded areas; and we observe that an intelligent choice of $\epsilon$ can make a huge difference.
@@
Dataset Creation. First, we obtain a set of $n$ correctly classified (genuine) inputs from the test data, $X_{g}=\{\vec{x}^{g}_{1},\ldots,\vec{x}^{g}_{n}\}$. From this set, we construct a set $X_{a}=\{\vec{x}^{a}_{1},\ldots,\vec{x}^{a}_{m}\}$ of adversarial inputs, obtained from the points in $X_{g}$ by applying state-of-the-art attacks, such as PGD [24, 26]. We also consider a set of $\ell$ potential $\epsilon$ values, $E=\{\epsilon_{1},\ldots,\epsilon_{\ell}\}$, from among which we will select the $\epsilon$ value to be used during inference.
@@
Next, for each $\epsilon\in E$ and for each input point $\vec{x}\in X=X_{g}\cup X_{a}$, we sample $k$ perturbed inputs $\vec{x}^{\prime}_{1}\ldots\vec{x}^{\prime}_{k}$ from the $L_{\infty}$ $\epsilon$-region around $\vec{x}$ [21, 22]. We then evaluate $N$ on $\vec{x}$ and on its perturbations $\vec{x}^{\prime}_{1}\ldots\vec{x}^{\prime}_{k}$, obtaining the outputs $N(\vec{x})$ for $\vec{x}$ and $N(\vec{x}^{\prime}_{1})\ldots N(\vec{x}^{\prime}_{n})$ for the perturbations. We count the number of hits, $h^{\vec{x}}_{\epsilon}=|\{i\ |\operatorname*{arg\,max}{N(\vec{x}^{\prime}_{i})}=\operatorname*{arg\,max}{N(\vec{x})}\}|$, and store it for later use.
@@
We note that instead of using the measure of hits as a predictor for whether the input being examined is adversarial, one could try to discover the distribution of adversarial inputs, and then attempt to decide whether an input is likely to belong to that distribution. However, determining the distribution of adversarial inputs is difficult, and could not be achieved, empirically, using state-of-the-art statistical tools [27].
4.3 Maximal-Recall-Oriented Calibration
@@
Using the collected data for various $\epsilon$ values, we present Alg. 1 for selecting an optimal value of $\epsilon$, and its corresponding threshold $\mathcal{T}$. The sought-after $\epsilon$ is optimal in the sense that, when used in the inference phase, it makes the fewest classification mistakes. The selection algorithm is brute-force: we consider each $\langle\epsilon,\mathcal{T}\rangle$ pair in turn, compute its success rate, and then pick the best pair.
@@
The algorithm receives as input the set of potential $\epsilon$ values $E$; the sets of genuine and adversarial points, $X_{g}$ and $X_{a}$, respectively; the measured number of hits, $h^{\vec{x}}_{\epsilon}$, for each input point $\vec{x}$ and each $\epsilon$; and also the hyper-parameter $w_{g}$, discussed later. The algorithm’s outputs are the selected $\epsilon$ value, denoted $\epsilon^{*}$, and its corresponding threshold of hits, $\mathcal{T}$, to be used in the inference phase.
@@
Line 2 constitutes the main loop of the algorithm, iterating over all potential values of $\epsilon$ in order to pick the best one. For each such $\epsilon$, we iterate over all possible threshold values in Line 3; since each $\langle\epsilon,\vec{x}\rangle$ pair was tested over $k$ perturbed inputs, the number of hits can be anywhere between 00 and $k$, and we seek the best choice. Line 4 computes the fraction of genuine inputs for which the threshold was reached, whereas Line 5 computes the fraction of adversarial inputs for which the threshold was not reached; these values are called the recall rates, and are denoted $r_{g}$ and $r_{a}$, respectively. Then, Line 6 computes a real-valued score for this candidate $\epsilon$. The score is a weighted average between $r_{g}$ and $r_{a}$, where the hyper-parameter $w_{g}$ is used in prioritizing between a more conservative classification of inputs ($w_{g}$ is lower, leading to fewer adversarial inputs mistakenly classified as genuine), or a less conservative one. The best score, and its corresponding $\epsilon$ and threshold, are stored in Line 8; and the selected $\epsilon$ and threshold are returned in Line 12.
Note. Alg. 1 computes a single $\epsilon$ value to be used during inference. However, in practice we observed that the approach is more successful when different $\epsilon$ values are used for the different output classes (see Fig. 1). E.g., during inference, if the input at hand is classified as “dog”, one $\epsilon$ value is used; whereas if it classified as “cat”, another is selected. To make this adjustment, Alg. 1 needs to be run once for each output class, and the sets $X_{a}$ and $X_{g}$ need to be partitioned by output class, as well. For brevity, we omit these adjustments here, as well as in subsequent algorithms.
4.4 Maximal-Precision-Oriented Calibration
@@
Alg. 1 is geared for recall; i.e., during inference, it classifies every input as either genuine (if the number of hits exceeds the threshold), or as suspected as adversarial. However, in our studies we observed that often, a division into three classes was more appropriate: in many cases, inputs encountered during inference would either demonstrate a very high number of hits, in which case they are likely genuine; a very low number of hits, in which case they are likely adversarial; or a medium number of hits, in which case it is unclear whether they are adversarial or not. Whereas in Alg. 1 a human would have to inspect all inputs suspected as adversarial, a more subtle division into there classes could allow the human to only inspect those inputs that are in the third, “unknown” category. See Fig. 2 for an illustration.
To support this variant, we present Alg. 2, which is geared for increased precision. It is similar to Alg. 1, but instead of selecting an $\epsilon$ and returning a single threshold $\mathcal{T}$, it now returns a couple of thresholds — $\mathcal{T}_{g}$ that indicates a lower bar of hits for genuine inputs, and $\mathcal{T}_{a}$ that indicates an upper bar of hits for adversarial inputs. When this algorithm is used, the final inference step is changed into:
@@
We observe that when using this three-output scheme, there is an inherent tension between precision and recall. For example, remembering that genuine inputs tend to have a high number of hits and adversarial inputs a low number, one could set $\mathcal{T}_{g}$ to be very high and $\mathcal{T}_{a}$ to be very low. This would result in excellent precision (i.e., the algorithm would make only few mistakes), but many inputs would fall in the “unknown” category, causing poor recall. In terms of Fig. 2, this would result in a very wide yellow strip. Alternatively, one could set these two thresholds to be near-identical (resulting in a narrow yellow strip), which would simply mimic Alg. 1, with its excellent recall but sub-optimal precision. To resolve this, we allow the user to specify a-priori the desired precision, by setting minimal values $p_{g}^{\min}$ and $p_{a}^{\min}$ for the precision over genuine and adversarial inputs, respectively. Our algorithm then automatically selects the $\epsilon$ and threshold values that meet these requirements, and which achieve the best recall among all candidates.
More concretely, the inputs to Alg. 2 are the same as those to Alg. 1, plus the two threshold parameters, $p_{g}^{\min}$ and $p_{a}^{\min}$. For each candidate $\epsilon$, the algorithm iterates over all possible thresholds (Line 4), from high to low, each time computing the recall rates (Lines 56) and the precision rate over genuine inputs (Line 7). If the precision rate meets the specified threshold, the threshold is stored as $t_{1}$ in Line 9. At the end of this loop, the algorithm has identified the lowest threshold for genuine inputs that meets the requirements. Next, the algorithm begins a symmetrical process for the threshold for adversarial inputs, this time iterating from low to high (Line 14), until it identifies the greatest threshold that satisfies the requirements ($t_{2}$, Line 19). Once both thresholds are discovered, the algorithm uses the recall ratios to compute a score (Line 24) — in order to compare the various $\epsilon$ values. The best $\epsilon$ and its thresholds are stored in Line 25, and are finally returned in Line 29.
@@ 5 Evaluation
Implementation. In order to evaluate our method, we created a proof-of-concept implementation of Algs. 1 and 2 in a tool called DEM, which is available online and will be made publicly available with the final version of this paper [16]. Our tool is written in Python, and supports DNNs in the common PyTorch format.
@@
Baseline. In our experiments we compared the two algorithms within DEM (recall-oriented and precision-oriented), and also compared DEM’s recall-oriented algorithm to the Local Intrinsic Dimension (LID) method [23], which is the state-of-the-art detecting adversarial examples. Like DEM, LID has a calibration phase, in which the network is evaluated on multiple genuine and adversarial inputs. LID uses these evaluations to examine the assignments to the various neurons of the DNN, and then trains a regression model to predict, during inference and based on these values, whether an input is adversarial. Since LID was not originally designed for categorical data, we trained it using two separate methods: (i) using examples from all classes, as originally suggested [23]; and (ii) training ten LID detectors, one for each output class. The second approach improved LID’s performance slightly, and we used it in the experiments described next.
@@
Benchmarks. We used VGG16 [29] and Resnet [13] DNN models, trained on the CIFAR10 dataset [18]. The Resnet and VGG models achieved $88.35\%$ and $83.75\%$ accuracy, respectively. We then took the CIFAR10 test set, removed from it any inputs that were misclassified by the trained models, and split the remaining inputs into two sets: a calibration set with $80\%$ of the inputs, and an evaluation set with the remaining $20\%$. The adversarial inputs for the calibration and evaluation phases were obtained using PGD [24], with a maximal modification distance of $0.005$ (i.e., $0.5\%$).
@@
In the preliminary calibration phase, we created $1000$ perturbed inputs around each genuine and adversarial example. We empirically observed that $1000$ samples was a suitable choice sufficient — using additional inputs did not change our algorithm’s outputs significantly, whereas using significantly fewer inputs led to a less-accurate calibration. For both calibration methods, we set the genuine recall hyperparameter to $w_{g}=0.3$. This conservative choice prioritizes avoiding adversarial examples wrongly misclassified as genuine. For the set $E$ of candidate $\epsilon$ values, we arbitrarily used distance values ranging from $0.001$ to $0.15$, with $0.001$ resolution. While RoMA and gRoMA used an $\epsilon$ value of $0.04$, we decided to use a greater value to ensure we do not miss a better calibration value. We note again that large values of $\epsilon$ reduce the number of hits for genuine inputs, because the region includes legitimate inputs that have true different classifications; and we empirically found that a distance greater than $0.15$ typically causes this number to drop too steeply.
@@
We conducted our evaluation on a standard laptop, equipped with an AMD Ryzen 9 6900HX CPU, an NVIDIA GeForce RTX 3070 GPU, and 16GB of RAM. In order to create a virtual running environment, we used Python 3.11 and PyTorch 2.11. The calibration and data preparation for each algorithm took less than four hours, per class. Each sample was analyzed in less than 0.003 seconds. The code for the creation of adversarial examples, calibration, and evaluation is available online [16].
5.1 Evaluating the Recall-Oriented Calibration Algorithm
@@
Table 1 provides a comparison of our recall-oriented calibration algorithm with LID [23]. It demonstrates that the proposed method is highly effective at distinguishing between genuine and adversarial inputs. Compared to LID, we observe that LID identifies genuine examples with a slightly better recall; however, DEM is significantly better at identifying adversarial inputs significantly, for all classes. This is a favorable result, as the mistake of classifying an adversarial input as a genuine one is considered to be much more serious than the second kind of mistake — classifying a genuine input as an adversarial.
@@
Table 1: Evaluating the recall-oriented calibration algorithm, in comparison to LID. The evaluation includes the Resnet (marked as RES) and VGG (marked as VGG) models. The two parts of the table present genuine and adversarial recall: the number of samples, the recall achieved by DEM, and the recall achieved by LID.
@@
#Samples RES|VGG DEM LID #Samples RES|VGG DEM LID
Airplane 180|177 63%|76% 81%|88% 393|125 92%|86% 54%|30%
Automotive 190|185 65%|65% 80%|91% 290|80 94%|87% 43%|52%
Bird 169|160 79%|67% 80%|84% 441|159 91%|91% 44%|35%
Cat 149|146 51%|50% 74%|85% 547|246 91%|89% 52%|31%
Deer 179|176 80%|79% 76%|89% 526|186 93%|97% 47%|37%
Dog 164|135 56%|60% 74%|87% 465|183 91%|92% 44%|30%
Frog 183|174 99%|91% 76%|84% 452|138 99%|97% 54%|34%
Horse 183|174 65%|74% 84%|88% 404|125 88%|87% 44%|38%
Ship 188|185 88%|91% 79%|92% 260|64 96%|92% 57%|56%
Truck 185|184 86%|80% 83%|87% 331|100 95%|95% 45%|44%
Examining the result, we observe the notable effectiveness of DEM when applied to the specific classes of the Resnet model. For instance, note the Frog class, where detection rates consistently surpass 90% across numerous samples. Nevertheless, some of the other classes demonstrate lesser effectiveness, and consequently these results, as a whole, do not meet the rigorous standards of aerospace certification.
5.2 Evaluating the Precision-Oriented Calibration Algorithm
We conducted another, similar experiment, this time using precision-oriented calibration, with precision thresholds set to $85\%$ for genuine examples on Resnet and VGG, and $80\%$ for adversarial inputs with both models. The $w_{g}$ hyperparameter remains the same as in the recall evaluation. Table 2 depicts the results. The recall results reported in the table appear alongside their corresponding precision results, for the same values of $\epsilon$.
Except for one case, all recall values were above $50\%$, and most of them were over $70\%$. As expected, our second algorithm produces a detector with the required precision value in most cases.
Table 2: Evaluation the precision-oriented calibration algorithm. Resnet results are marked as RES, and VGG results as VGG.

### v3

@@ IV-B Dataset Preparation and Calibration
Figure 1: An illustration of CIFAR-10 classifier performance. Each plot corresponds to a single output category. The $Y$-axes in the plots show the average number of hits (bold line) and standard deviation (faded area), for $k=1000$ perturbations around genuine (blue) and adversarial (orange) inputs; whereas the $X$-axes represent different values of $\epsilon$ (in percents). The goal is to have as small overlap as possible between the two distributions, observing the significance of $\epsilon$ and the variance of the distributions between different categories.
@@
Dataset Creation. First, we obtain a set of $n$ correctly classified (genuine) inputs from the test data, $X_{g}=\{\vec{x}^{g}_{1},\ldots,\vec{x}^{g}_{n}\}$. From this set, we construct a set $X_{a}=\{\vec{x}^{a}_{1},\ldots,\vec{x}^{a}_{n_{a}}\}$ of adversarial inputs, obtained from the points in $X_{g}$ by applying state-of-the-art attacks, such as PGD [37, 38]. We also consider a set of $m$ potential $\epsilon$ values, $E=\{\epsilon_{1},\ldots,\epsilon_{m}\}$, from among which we will select the $\epsilon$ value to be used during inference.
@@
Next, for each $\epsilon\in E$ and for each input point $\vec{x}\in X=X_{g}\cup X_{a}$, we sample $k$ perturbed inputs $\vec{x}^{\prime}_{1}\ldots\vec{x}^{\prime}_{k}$ from the $L_{\infty}$ $\epsilon$-region around $\vec{x}$ [35, 13]. We then evaluate $N$ on $\vec{x}$ and on its perturbations $\vec{x}^{\prime}_{1}\ldots\vec{x}^{\prime}_{k}$, obtaining the outputs $N(\vec{x})$ for $\vec{x}$ and $N(\vec{x}^{\prime}_{1})\ldots N(\vec{x}^{\prime}_{k})$ for the perturbations. We count the number of hits, $h^{\vec{x}}_{\epsilon}=|\{i\ |\operatorname*{arg\,max}{N(\vec{x}^{\prime}_{i})}=\operatorname*{arg\,max}{N(\vec{x})}\}|$, and store it for later use.
@@
We note that instead of using the measure of hits as a predictor for whether the input being examined is adversarial, one could try to discover the distribution of adversarial inputs, and then attempt to decide whether an input is likely to belong to that distribution. However, determining the distribution of adversarial inputs is difficult, and presently cannot be achieved using state-of-the-art statistical tools [39].
IV-C Maximal-Recall-Oriented Calibration
@@
Using the collected data for various $\epsilon$ values, we present Alg. 1 for selecting an optimal $\epsilon$ value, and its corresponding threshold $\mathcal{T}$. The sought-after $\epsilon$ is optimal in the sense that, when used in the inference phase, it makes the fewest classification mistakes. The selection algorithm is brute-force: we consider each $\langle\epsilon,\mathcal{T}\rangle$ pair in turn, compute its success rate, and then pick the most successful pair.
@@
The algorithm receives as input the set of potential $\epsilon$ values $E$; the sets of genuine and adversarial points, $X_{g}$ and $X_{a}$, respectively; the measured number of hits, $h^{\vec{x}}_{\epsilon}$, for each input point $\vec{x}$ and each $\epsilon$; and also the hyperparameter $w_{g}$, discussed later. The algorithm’s outputs are the selected $\epsilon$ value, denoted $\epsilon^{*}$, and its corresponding threshold of hits, $\mathcal{T}$, to be used during the inference phase.
@@
Line 2 is the algorithm’s main loop, iterating over all potential $\epsilon$ values in order to pick the optimal one. For each such $\epsilon$, we iterate over all possible threshold values in Line 3; since each pair $\langle\epsilon,\vec{x}\rangle$ was tested over $k$ perturbed inputs, the number of hits may vary between 00 and $k$, and we seek the best choice. Line 4 calculates the fraction of genuine inputs for which the threshold was reached, whereas Line 5 computes the fraction of adversarial inputs for which the threshold was not reached; these values are called the recall rates, and are denoted $r_{g}$ and $r_{a}$, respectively. Then, Line 6 computes a real-valued score for this candidate $\epsilon$. The score is a weighted average between $r_{g}$ and $r_{a}$, where the hyperparameter $w_{g}$ is used in prioritizing between a more conservative classification of inputs ($w_{g}$ is lower, leading to fewer adversarial inputs mistakenly classified as genuine), or a less conservative one. The best score, and its corresponding $\epsilon$ and threshold, are stored in Line 8; and the selected $\epsilon$ and threshold are returned in Line 12.
Note. Alg. 1 computes a single $\epsilon$ value to be used during inference. However, in practice we used different $\epsilon$ value for each of the different output classes (see Fig. 1). E.g., during inference, if the input at hand is classified as “Dog”, one $\epsilon$ value is used; whereas if it is classified as “Cat”, another is selected. To make this adjustment, Alg. 1 needs to be run once for each output class, and the sets $X_{a}$ and $X_{g}$ need to be partitioned by output class, as well. For brevity, we omit these adjustments here, as well as in subsequent algorithms.
IV-D Maximal-Precision-Oriented Calibration
@@
Alg. 1 is geared for recall; i.e., during inference, it classifies every input as either genuine (if the number of hits exceeds the threshold), or as adversarial. However, in our studies we observed that often, a division into three classes was more appropriate: in many cases, inputs encountered during inference would either demonstrate a very high number of hits, in which case they are likely genuine; a very low number of hits, in which case they are likely adversarial; or a medium number of hits, in which case it is unclear whether they were adversarial or not. While in Alg. 1 a human would have to inspect all inputs classified as adversarial, a more subtle division into three classes could allow the human to only inspect those inputs in the third, “unknown” category. See Fig. 2 for an illustration.
To support this variant, we present Alg. 2, which is geared towards increased precision. It is similar to Alg. 1, but instead of selecting an $\epsilon$ and returning a single threshold $\mathcal{T}$, it now returns a couple of thresholds — $\mathcal{T}_{g}$, which indicates a lower bar of hits for genuine inputs, and $\mathcal{T}_{a}$, which indicates an upper bar of hits for adversarial inputs. When this algorithm is used, the final inference step is changed to:
@@
We observe that this three-output scheme has an inherent tension between precision and recall. For example, remembering that genuine inputs tend to have a high number of hits and adversarial inputs a lower number, one could set $\mathcal{T}_{g}$ to be very high and $\mathcal{T}_{a}$ to be very low. This would result in excellent precision (i.e., the algorithm would make only a few mistakes), but many inputs would fall into the “unknown” category, causing poor recall. In terms of Fig. 2, this would result in a very wide yellow strip. Alternatively, one could set these two thresholds to be near-identical (resulting in a narrow yellow strip), which would simply mimic Alg. 1, with its excellent recall but sub-optimal precision. To resolve this, we allow the user to specify a-priori the desired precision, by setting minimal values $p_{g}^{\min}$ and $p_{a}^{\min}$ for the precision over genuine and adversarial inputs, respectively. Our algorithm then automatically selects the $\epsilon$ and threshold values that meet these requirements, and which achieve the maximum recall among all candidates.
More concretely, the inputs to Alg. 2 are the same as those to Alg. 1, plus the two threshold parameters, $p_{g}^{\min}$ and $p_{a}^{\min}$. For each candidate $\epsilon$, the algorithm iterates over all possible thresholds (Line 4), from high to low, each time computing the recall rates (Lines 56) and the precision rate over genuine inputs (Line 7). If the precision rate meets the specified threshold, the threshold is stored as $t_{1}$ in Line 9. At the end of this loop, the algorithm has identified the lowest threshold for genuine inputs that meet the requirements. Next, the algorithm begins a symmetrical process for the threshold for adversarial inputs, this time iterating from low to high (Line 14), until it identifies the greatest threshold that satisfies the requirements ($t_{2}$, Line 19). Once both thresholds are discovered, the algorithm uses the recall ratios to compute a score (Line 24) — in order to compare the various $\epsilon$ values. The most optimal $\epsilon$ and its thresholds are stored in Line 25, and are finally returned in Line 29.
@@ V Evaluation
Implementation. In order to evaluate our method, we created a proof-of-concept implementation of Algs. 1 and 2 in a tool called DEM, which is available online [40]. Our tool is written in Python, and supports DNNs in the common PyTorch format.
@@
Baseline. In our experiments we compared the two algorithms within DEM (recall-oriented and precision-oriented), and also compared DEM’s recall-oriented algorithm to the Local Intrinsic Dimension (LID) method [31], which is the state of the art in detecting adversarial examples. Like DEM, LID has a calibration phase, in which the network is evaluated on multiple genuine and adversarial inputs. LID uses these evaluations to examine the assignments to the various neurons of the DNN, and then trains a regression model to predict, during inference and based on these values, whether an input is adversarial.
Since LID was not originally designed for categorical data, we trained it using two separate methods: (i) using examples from all classes, as originally suggested [31]; and (ii) training ten LID detectors, one for each output class. The second approach improved LID’s performance slightly, and we used it as benchmark in the experiments described next.
@@
Benchmarks. We used VGG16 [14] and Resnet [15] DNN models, trained on the CIFAR10 dataset [41]. The Resnet and VGG models achieved $88.35\%$ and $83.75\%$ accuracy, respectively. We then took the CIFAR10 test set, and removed from it any inputs that were misclassified by the trained models. We split the remaining inputs into two sets: a calibration set with $80\%$ of the inputs, and an evaluation set with the remaining $20\%$. The adversarial inputs for the calibration and evaluation phases were obtained using PGD [37], with a maximal modification distance of $0.005$ (i.e., $0.5\%$).
@@
In the preliminary calibration phase, we created $1000$ perturbed inputs around each genuine and adversarial input. We empirically observed that $1000$ samples was a suitable choice — adding additional samples did not change our algorithm’s performance significantly, whereas using significantly fewer samples led to degraded performance. For both calibration methods, we set the genuine recall hyperparameter to $w_{g}=0.3$. This conservative choice prioritizes avoiding adversarial inputs wrongly misclassified as genuine. For the set $E$ of candidate $\epsilon$ values, we arbitrarily chose distance values ranging from $0.001$ to $0.15$, with $0.001$ resolution. While RoMA and gRoMA used an $\epsilon$ value of $0.04$ [35, 13], we decided to use a higher value to ensure we do not miss a better calibration value.
We note that large values of $\epsilon$ reduce the number of hits for genuine inputs, because the region includes legitimate inputs with true different classifications; we empirically found that a distance greater than $0.15$ typically causes this number to drop steeply.
@@
We conducted our evaluation on a standard laptop, equipped with an AMD Ryzen 9 6900HX CPU, an NVIDIA GeForce RTX 3070 GPU, and 16GB of RAM. In order to create a virtual running environment, we used Python 3.11 and PyTorch 2.11. The calibration and data preparation for each algorithm took less than four hours per class. Each sample was analyzed in less than 0.003 seconds. The code for the creation of adversarial examples, calibration, and evaluation is available online [40].
V-A Evaluating the Recall-Oriented Calibration Algorithm
@@
Table V-A provides a comparison of our recall-oriented calibration algorithm with LID [31]. It demonstrates that the proposed method is highly effective at distinguishing between genuine and adversarial inputs. Compared to LID, we observe that LID identifies genuine examples with a slightly better recall; however, DEM is significantly better at identifying adversarial inputs for all classes. This is a favorable result, as the mistake of classifying an adversarial input as a genuine one is considered to be much more serious, in the use-cases we consider, than the second kind of mistake — classifying a genuine input as an adversarial.
@@
TABLE I: Evaluating the recall-oriented calibration algorithm, in comparison to LID. The evaluation includes the Resnet (marked as RES) and VGG (marked as VGG) models. The two parts of the table present genuine and adversarial recall: the number of samples, the recall achieved by DEM, and the recall achieved by LID.
@@
#Samples RES, VGG DEM LID #Samples RES, VGG DEM LID
Airplane 180,177 63%,76% 81%,88% 393,125 92%,86% 54%,30%
Automotive 190,185 65%,65% 80%,91% 290,80 94%,87% 43%,52%
Bird 169,160 79%,67% 80%,84% 441,159 91%,91% 44%,35%
Cat 149,146 51%,50% 74%,85% 547,246 91%,89% 52%,31%
Deer 179,176 80%,79% 76%,89% 526,186 93%,97% 47%,37%
Dog 164,135 56%,60% 74%,87% 465,183 91%,92% 44%,30%
Frog 183,174 99%,91% 76%,84% 452,138 99%,97% 54%,34%
Horse 183,174 65%,74% 84%,88% 404,125 88%,87% 44%,38%
Ship 188,185 88%,91% 79%,92% 260,64 96%,92% 57%,56%
Truck 185,184 86%,80% 83%,87% 331,100 95%,95% 45%,44%
Examining the results, we observe the notable effectiveness of DEM when applied to the specific classes of the Resnet model. For instance, note the Frog class, where detection rates consistently surpass 90% across numerous samples. Nevertheless, some of the other classes demonstrate lesser effectiveness, and consequently these results, as a whole, do not meet the aerospace certification standards.
V-B Evaluating the Precision-Oriented Calibration Algorithm
We conducted another similar experiment, this time using precision-oriented calibration, with precision thresholds set to $85\%$ for genuine examples on Resnet and VGG, and $80\%$ for adversarial inputs with both models. The $w_{g}$ hyperparameter remained the same as in the recall evaluation. Table V-B depicts the results. The recall results reported therein appear alongside their corresponding precision results, for the same $\epsilon$ values. Except for one case, all recall values were above $50\%$, and most of them were over $70\%$. That second algorithm produces a detector with the required precision value in most cases.
TABLE II: Evaluation of the precision-oriented calibration algorithm. Resnet results are marked as RES, and VGG results as VGG.